OpenStudy (shaniehh):

PLEASE HELP!!!! the midpoint of the segment with endpoints at (-1,y) and (y,16) is (3,4). what is the value of y

OpenStudy (shaniehh):

@jim_thompson5910

OpenStudy (shaniehh):

@ganeshie8

OpenStudy (shaniehh):

@Hero @aaronq

jimthompson5910 (jim_thompson5910):

to find the midpoint, you average the corresponding coordinates

jimthompson5910 (jim_thompson5910):

what that means is that you add up the corresponding coordinates, then divide by 2

jimthompson5910 (jim_thompson5910):

so say we had these two points A(2, 3) B(6, 9) We would average the x coordinates (x1+x2)/2 = (2+6)/2 = 8/2 = 4 then average the y coordinates (y1+y2)/2 = (3+9)/2 = 12/2 = 6 So the midpoint is (4,6). This is just an example of course and not the answer.

OpenStudy (shaniehh):

can u please show me how to do the problem above its for my little brother im embarrassed I can help him

OpenStudy (shaniehh):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

here are the two points (-1,y) and (y,16)

jimthompson5910 (jim_thompson5910):

what are the x coordinates of each point?

OpenStudy (shaniehh):

ok so -1 -y & y-16

jimthompson5910 (jim_thompson5910):

-1 and y are the x coordinates of (-1,y) and (y,16)

jimthompson5910 (jim_thompson5910):

add those two x coordinates up and divide by 2 (-1+y)/2

jimthompson5910 (jim_thompson5910):

the midpoint is (3,4)

jimthompson5910 (jim_thompson5910):

the x coordinate of the midpoint is 3 that means (-1+y)/2 = 3

OpenStudy (shaniehh):

wait is that the answer or do i have to make any more steps

jimthompson5910 (jim_thompson5910):

you now have to solve for y